Aspartame is an artificial sweetener used in making desserts,cakes,soft drinks,and others.its molecular formula is C14H18N2O3.determine the percentage composition
1
Expert's answer
2020-12-10T07:48:46-0500

The correct formula of Aspartame is C14H18N2O5

The molecular weight - 294 g/mol

Answer:

Percentage of C: (12 * 14) / 294 = 57.15%

Percentage of H: (1 * 18) / 294 = 6.12 %

Percentage of N: (14 * 2) / 294 = 9.52 %

Percentage of O: (16 * 5) / 294 = 27.21 %